What is the Na+ concentration in each of the following solutions:

3.05 M sodium sulfate: _______M

1.18 M sodium carbonate: ______M

0.685 M sodium bicarbonate: ______ M

(b) What is the concentration of a lithium carbonate solution that is 0.695 M in Li+? _______ M
Answer #1

a)

1)

Given:

[Na2SO4] = 3.05 M

So,

[Na+] = 2*[Na2SO4]

= 2*3.05 M

= 6.10 M

Answer: 6.10 M

2)

Given:

[Na2CO3] = 1.18 M

So,

[Na+] = 2*[Na2CO3]

= 2*1.18 M

= 2.36 M

Answer: 2.36 M

3)

Given:

[NaHCO3] = 0.685 M

So,

[Na+] = 1*[NaHCO3]

= 0.685 M

Answer: 0.685 M

b)

Given:

[Li+] = 0.695 M

So,

[Li2CO3] = [Li+]/2

= 0.695/2

= 0.3475 M

Answer: 0.348 M